# Distributed Maximum Flow in Planar Graphs

### v3

@@ Abstract
Given a directed planar graph $G$ with positive and negative edge-lengths and hop-diameter $D$, our key result is an $\tilde{O}(D^{2})$-round algorithm [^1] The $\tilde{O}(\cdot)$ notation is used to omit $\text{poly}\log n$ factors. for Single Source Shortest Paths on $G^{*}$. This algorithm implies an $\tilde{O}(D^{2})$-round algorithm for Maximum $st$-Flow in $G$, and an $\tilde{O}(D^{2})$-round algorithm for directed Global Minimum Cut in $G$. Prior to our work, no $\tilde{O}(\text{poly}(D))$-round algorithms were known for these problems. When $G$ is undirected, we further obtain a near optimal $\tilde{O}(D)$-round algorithm for computing the weighted girth of $G$, and a $D\cdot n^{o(1)}$-rounds $(1-\epsilon)$-approximation algorithm for Maximum $st$-Flow in $G$ when $s$ and $t$ lie on the same face.
The main challenges in our work are that $G^{*}$ is not the communication graph (e.g., a vertex of $G$ is mapped to multiple vertices of $G^{*}$), and that the diameter of $G^{*}$ can be much larger than $D$ (i.e., possibly by a linear factor). We overcome these challenges by carefully defining and maintaining subgraphs of the dual graph $G^{*}$ while applying the recursive decomposition on the primal graph $G$. The main technical difficulty, is that along the recursive decomposition, a face of $G$ gets shattered into (disconnected) components yet we still need to treat it as a dual node.
@@ 1 Introduction
Distributed algorithms for network optimization problems have a long and rich history. These problems are commonly studied under the $\mathsf{CONGEST}$ model [36] where the network is abstracted as an $n$-vertex graph $G=(V,E)$ with hop-diameter $D$; communications occur in synchronous rounds, and per round, $O(\log n)$ bits can be sent along each edge. A sequence of breakthrough results provided $\tilde{O}(D+\sqrt{n})$-round algorithms for fundamental graph problems, such as minimum spanning tree (MST) [9], approximate shortest-paths [33], minimum cuts [5], and approximate flow [16]. For general graphs, $\tilde{O}(D+\sqrt{n})$ rounds for solving the above mentioned problems is known to be near optimal, existentially [39].
@@
An additional key tool in working with planar graphs, starting with the seminal work of Lipton and Tarjan [27], is that of a planar separator path: a path whose removal from the graph leaves connected components that are a constant factor smaller. Ghaffari and Parter [17] presented a $\tilde{O}(D)$-round randomized algorithm for computing a cycle separator of size $O(D)$ which consists of a separator path plus one additional edge (that is possibly a virtual edge that is not in $G$). By now, planar separators are a key ingredient in a collection of $\tilde{O}(\text{poly}(D))$-round solutions for problems such as DFS [17], distance computation [26], and reachability [35]. An important aspect of the planar separator algorithm of [17] is that it employs a computation on the dual graph, by communicating over the primal graph.
Primal maximum flow via dual SSSP. Our goal in this paper is to expand the algorithmic toolkit for performing computation on the dual graph. This allows us to exploit the profound algorithmic duality in planar graphs, in which solving a problem $A$ in the dual graph provides a solution for problem $B$ in the primal graph. Within this context, our focus is on the Maximum $st$-Flow problem (in directed planar graphs with edge capacities). The Maximum $st$-Flow problem is arguably one of the most classical problems in theoretical computer science, extensively studied since the 50’s, and still admitting breakthrough results in the sequential setting, such as the recent almost linear time algorithm by Chen, Kyng, Liu, Peng, Gutenberg and Sachdeva [2]. Despite persistent attempts over the years, our understanding of the distributed complexity of this problem is still quite lacking. For general undirected $n$-vertex graphs, there is a $(1+o(1))$-approximation algorithm that runs in $(\sqrt{n}+D)n^{o(1)}$ rounds, by Ghaffari, Karrenbauer, Kuhn, Lenzen and Patt-Shamir [16]. For directed $n$-vertex planar graphs, a $D\cdot n^{1/2+o(1)}$-round exact algorithm has been given by de Vos [4]. No better tradeoffs are known for undirected planar graphs. In lack of any $\tilde{O}(\text{poly}(D))$-round maximum $st$-flow algorithm for directed planar graphs (not even when allowing approximation) we ask:
@@
In directed planar graphs with integral edge-capacities, it is known from the 80’s [45] that the maximum $st$-flow can be found by solving $\log\lambda$ instances of Single Source Shortest Paths (SSSP) with positive and negative edge-lengths on the dual graph $G^{*}$, where $\lambda$ is the maximum $st$-flow value. We answer 1.1 in the affirmative by designing a $\tilde{O}(D^{2})$-round SSSP algorithm on the dual graph $G^{*}$. Our algorithm works w.h.p. [^2] W.h.p. stands for a probability of $1-1/{n^{c}}$ for an arbitrary fixed constant $c>0$., it works in the most general setting (i.e. when $G^{*}$ is directed and has positive and negative integral edge-lengths) and it matches the fastest known exact SSSP algorithm in the primal graph. We show:
@@
There is a randomized distributed algorithm that given an $n$-vertex directed planar network $G$ with hop-diameter $D$ and integral edge-capacities, and two vertices $s,t$, computes the maximum $st$-flow value and assignment w.h.p. in $\tilde{O}(D^{2})$ rounds.
No prior $\tilde{O}(\text{poly}(D))$ algorithm has been known for this problem, not even when allowing a constant approximation. We further improve the running time to $D\cdot n^{o(1)}$ rounds while introducing a $(1-\epsilon)$ approximation, provided that $G$ is undirected and that $s$ and $t$ both lie on the same face with respect to the given planar embedding:
@@
There is a randomized distributed algorithm that given an $n$-vertex undirected planar network $G$ with hop-diameter $D$ and integral edge-capacities, and two vertices $s,t$ lying on the same face, computes a $(1-\epsilon)$-approximation of the maximum $st$-flow value and a corresponding assignment in $D\cdot n^{o(1)}$ rounds w.h.p.
@@
Directed global minimum cut. The directed global minimum weight cut problem is given a weighted directed graph and asks to find a bisection $(S,V\setminus S)$ of the vertex set, such that, the set of edges leaving $S$ to $V\setminus S$ is of minimal total weight. To our knowledge, this problem was not studied in the $\mathsf{CONGEST}$ model and is open even for approximations, as opposed to the (directed or undirected) minimum $st$-cut problem [4, 34] and the undirected global minimum cut problem [18, 5, 14]. In light of efficient planar centralized results [31, 46] and motivated by the importance of the problem, we ask:
@@
Primal weighted girth via dual cuts. A distance parameter of considerable interest is the network girth. For unweighted graphs, the girth is the length of the shortest cycle in the graph. For weighted graphs, the girth is the cycle of minimal total edge weight. Distributed girth computation has been studied over the years mainly for general $n$-vertex unweighted graphs. Frischknecht, Holzer and Wattenhofer [8] provided an $\Omega(\sqrt{n})$-round lower bound for computing a $(2-\epsilon)$ approximation of the unweighted girth. The state-of-the-art upper bound for the unweighted girth problem is a $(2-\epsilon)$ approximation in $\tilde{O}(n^{2/3}+D)$ rounds, obtained by combining the works of Peleg, Roditty and Tal [37] and Holzer and Wattenhofer [21]. The weighted girth problem has been shown to admit a near-optimal lower bound of $\tilde{\Omega}(n)$ rounds in general graphs [22, 28]. Turning to planar graphs, Parter [35] devised a $\tilde{O}(D^{2})$ round algorithm for computing the weighted girth in directed planar graphs via SSSP computations. For undirected and unweighted planar graphs, the (unweighted) girth can be computed in $\tilde{O}(D)$ rounds by replacing the $\tilde{O}(D^{2})$-round SSSP algorithm by a $O(D)$-round BFS algorithm. In light of this gap, we ask:
@@ 2.1 Minor-Aggregation in the Dual
An important recent development in the field of distributed computing was a new model of computation, called the minor-aggregation model introduced by Zuzic ⓡ444ⓡ is used to denote that the authors’ ordering is randomized. Goranci ⓡ Ye ⓡ Haeupler ⓡ Sun [42], then extended by Ghaffari and Zuzic [18] to support working with virtual nodes added to the input graph. Recent state-of-art algorithms for various classical problems can be formulated in the minor-aggregation model (e.g., the exact min-cut algorithm of [18], and the undirected shortest paths approximation algorithms of [41, 42]). Motivated by the algorithmic power of this model, we provide an implementation of the minor aggregation model on the dual graph. As noted by [42], minor aggregations can be implemented by solving the (simpler) part-wise aggregation task, where one needs to compute an aggregate function in a collection of vertex-disjoint connected parts of the graph. The distributed planar separator algorithm of [17] implicitly implements a part-wise aggregation algorithm on the dual graph. Our contribution is in providing an explicit and generalized implementation of the dual part-wise aggregation algorithm and using it to implement the minor-aggregation model on the dual graph, see Section 4 for the details. We then use this algorithm for computing the exact minimum weighted cut in the dual graph $G^{*}$, which by duality provides a solution to the weighted girth problem in the primal graph $G$. We also use it to simulate the recent approximate SSSP by [42] on the dual graph, leading to our approximate max $st$-flow algorithm on the primal graph when $s$ and $t$ are on the same face. Since currently there are fast SSSP minor-aggregation algorithms only for undirected graphs, this approach leads to an approximate max $st$-flow algorithm in undirected planar graphs. To solve the more general version of the max $st$-flow problem, we need additional tools described next.

### v4

# Yaseen Abd-Elhaleem

@@ Abstract
[…]
The main challenges in our work are that $G^{*}$ is not the communication graph (e.g., a vertex of $G$ is mapped to multiple vertices of $G^{*}$), and that the diameter of $G^{*}$ can be much larger than $D$ (i.e., possibly by a linear factor). We overcome these challenges by carefully defining and maintaining subgraphs of the dual graph $G^{*}$ while applying the recursive decomposition on the primal graph $G$. The main technical difficulty is that, along the recursive decomposition, a face of $G$ gets shattered into (disconnected) components, yet we still need to treat it as a dual node.
@@ 1 Introduction
Distributed algorithms for network optimization problems have a long and rich history. These problems are commonly studied under the $\mathsf{CONGEST}$ model [37] where the network is abstracted as an $n$-vertex graph $G=(V,E)$ with hop-diameter $D$; communications occur in synchronous rounds. In each round, $O(\log n)$ bits can be sent along every edge. A sequence of breakthrough results provided $\tilde{O}(D+\sqrt{n})$-round algorithms for fundamental graph problems, such as minimum spanning tree (MST) [9], approximate shortest-paths [34], minimum cuts [5], and approximate flow [16]. For general graphs, $\tilde{O}(D+\sqrt{n})$ rounds for solving the above mentioned problems is known to be near optimal, existentially [40].
@@
An additional key tool in working with planar graphs, starting with the seminal work of Lipton and Tarjan [28], is that of a planar path separator: a path whose removal from the graph leaves connected components that are a constant factor smaller. Ghaffari and Parter [17] presented a $\tilde{O}(D)$-round randomized algorithm for computing a cycle separator of size $O(D)$ which consists of a separator path plus one additional edge (that is possibly a virtual edge that is not in $G$). Recently, a deterministic construction of planar separators was given by Jauregui, Montealegre and Rapaport [23]. By now, planar separators are a key ingredient in a collection of $\tilde{O}(\text{poly}(D))$-round solutions for problems such as DFS [17, 23], distance computation [27], and reachability [36]. An important aspect of the planar separator algorithm of [17] is that it employs a computation on the dual graph, by communicating over the primal graph.
Primal maximum flow via dual SSSP. Our goal in this paper is to expand the algorithmic toolkit for performing computations on the dual graph. This allows us to exploit the profound algorithmic duality in planar graphs, in which solving a problem $A$ in the dual graph provides a solution for problem $B$ in the primal graph. Within this context, our focus is on the Maximum $st$-Flow problem (in directed planar graphs with edge capacities). The Maximum $st$-Flow problem is arguably one of the most classical problems in theoretical computer science, extensively studied since the 50’s, and still admitting breakthrough results in the sequential setting, such as the recent almost linear time algorithm by Chen, Kyng, Liu, Peng, Gutenberg and Sachdeva [2]. Despite persistent attempts over the years, our understanding of the distributed complexity of this problem is still quite lacking. For general undirected $n$-vertex graphs, there is a $(1+o(1))$-approximation algorithm that runs in $(\sqrt{n}+D)n^{o(1)}$ rounds, by Ghaffari, Karrenbauer, Kuhn, Lenzen and Patt-Shamir [16]. For directed $n$-vertex planar graphs, a $D\cdot n^{1/2+o(1)}$-round exact algorithm has been given by de Vos [4]. No better tradeoffs are known for undirected planar graphs. In lack of any $\tilde{O}(\text{poly}(D))$-round maximum $st$-flow algorithm for directed planar graphs (not even when allowing approximation) we ask:
@@
In directed planar graphs with integral edge-capacities, it is known from the 80’s [46] that the maximum $st$-flow can be found by solving $\log\lambda$ instances of Single Source Shortest Paths (SSSP) with positive and negative edge-lengths on the dual graph $G^{*}$, where $\lambda$ is the maximum $st$-flow value. We answer 1.1 in the affirmative by designing a $\tilde{O}(D^{2})$-round SSSP algorithm on the dual graph $G^{*}$. Our algorithm works w.h.p. [^2] W.h.p. stands for a probability of $1-1/{n^{c}}$ for an arbitrary fixed constant $c>0$. in the most general setting (i.e. when $G^{*}$ is directed and has positive and negative integral edge-lengths), and matches the fastest known exact SSSP algorithm in the primal graph. We show:
@@
There is a randomized distributed algorithm that, given an $n$-vertex directed planar network $G$ with hop-diameter $D$ and integral edge-capacities, and two vertices $s,t$, computes the maximum $st$-flow value and assignment w.h.p. in $\tilde{O}(D^{2})$ rounds.
No prior $\tilde{O}(\text{poly}(D))$ algorithm has been known for this problem, not even when allowing a constant approximation. We further improve the running time to $D\cdot n^{o(1)}$ rounds while introducing a (1o(1))(1-o(1)) approximation, provided that $G$ is undirected and that $s$ and $t$ both lie on the same face with respect to the given planar embedding:
@@
There is a randomized distributed algorithm that given an $n$-vertex undirected planar network $G$ with hop-diameter $D$ and integral edge-capacities, and two vertices $s,t$ lying on the same face, computes a (1o(1))(1-o(1))-approximation of the maximum $st$-flow value and a corresponding assignment in $D\cdot n^{o(1)}$ rounds w.h.p..
@@
Directed global minimum cut. The directed global minimum weight cut problem is given a weighted directed graph and asks to find a bisection $(S,V\setminus S)$ of the vertex set, such that, the set of edges leaving $S$ to $V\setminus S$ is of minimal total weight. To our knowledge, this problem was not studied in the $\mathsf{CONGEST}$ model and is open even for approximations, as opposed to the (directed or undirected) minimum $st$-cut problem [4, 35] and the undirected global minimum cut problem [18, 5, 14]. In light of efficient centralized results for planar graphs [32, 47] and motivated by the importance of the problem, we ask:
@@
Primal weighted girth via dual cuts. A distance parameter of considerable interest is the network girth. For unweighted graphs, the girth is the length of the shortest cycle in the graph. For weighted graphs, the girth is the cycle of minimal total edge weight. Distributed girth computation has been studied over the years mainly for general $n$-vertex unweighted graphs. Frischknecht, Holzer and Wattenhofer [8] provided an $\Omega(\sqrt{n})$-round lower bound for computing a $(2-\epsilon)$ approximation of the unweighted girth. The state-of-the-art upper bound for the unweighted girth problem is a $(2-\epsilon)$ approximation in $\tilde{O}(n^{2/3}+D)$ rounds, obtained by combining the works of Peleg, Roditty and Tal [38], and Holzer and Wattenhofer [21]. The weighted girth problem has been shown to admit a near-optimal lower bound of $\tilde{\Omega}(n)$ rounds in general graphs [22, 29]. Turning to planar graphs, Parter [36] devised a $\tilde{O}(D^{2})$ round algorithm for computing the weighted girth in directed planar graphs via SSSP computations. For undirected and unweighted planar graphs, the (unweighted) girth can be computed in $\tilde{O}(D)$ rounds by replacing the $\tilde{O}(D^{2})$-round SSSP algorithm by a $O(D)$-round BFS algorithm. In light of this gap, we ask:
@@ 2.1 Minor-Aggregation in the Dual
An important recent development in the field of distributed computing was a new model of computation, called the minor-aggregation model introduced by Zuzic ⓡ444ⓡis used to denote that the authors’ ordering is randomized. Goranci ⓡ Ye ⓡ Haeupler ⓡ Sun [43], then extended by Ghaffari and Zuzic [18] to support working with virtual nodes added to the input graph. Recent state-of-art algorithms for various classical problems are being formulated in the minor-aggregation model (e.g., the exact min-cut algorithm of [18], and the undirected shortest paths approximation algorithms of [42, 43]). Motivated by the algorithmic power of this model, we provide an implementation of the minor aggregation model on the dual graph. As noted by [43], minor aggregations can be implemented by solving the (simpler) part-wise aggregation task, where one needs to compute an aggregate function in a collection of vertex-disjoint connected parts of the graph. The distributed planar separator algorithm of [17] implicitly implements a part-wise aggregation algorithm on the dual graph. Our contribution is in providing an explicit and generalized implementation of the dual part-wise aggregation algorithm and using it to implement the minor-aggregation model on the dual graph. Then, using this implementation of the minor-aggregation model along with planar duality yields a collection of algorithmic results for the primal graph. Next, we give a brief overview of the proof of the simulation theorem, where the full details appear in Section 4.
The minor-aggregation model. In the minor-aggregation model, there is a given graph $G$, where the vertices and edges of $G$ are computational units, and an algorithm works in synchronous rounds, where in each round we can either contract some of the edges, or compute an aggregate function on certain sets of disjoint vertices or edges. For the full details of the model see Definition 4.7.
Part-wise aggregation. As shown in [43], the minor-aggregation model is simulated in $\mathsf{CONGEST}$ by solving the part-wise aggregation (PA) problem. Briefly, the problem considers the setting where a partition $\{G_{i}\}_{i=1}^{N}$ of $G$ is given s.t. each part $G_{i}$ is a connected subgraph of $G$, in addition each vertex $v\in V$ initially has some input $x_{v}$. The objective is then, for each subgraph $G_{i}$ to compute an aggregate operator over all inputs $x_{v}$ of vertices $v\in G_{i}$. An aggregate operator is a function that allows to replace $\tilde{O}(1)$-bit strings by one $\tilde{O}(1)$-bit string. This is usually a simple commutative function such as taking a minimum or a sum.
Low-congestion shortcuts. In their influential work, Ghaffari and Haeupler [13] define low-congestion shortcuts as follows. Let $\{G_{i}\}_{i=1}^{N}$ be a partition of $G$ into vertex-disjoint connected subgraphs, a $(\alpha,\beta)$-shortcut, is a set of subgraphs $\{H_{i}\}_{i=1}^{N}$ of $G$ such that the diameter of each $G_{i}\cup H_{i}$ is at most $\beta$ and each edge of $G$ participates in at most $\alpha$ subgraphs $H_{i}$.
Using this notion, the part-wise aggregation task is solved easily by constructing low-congestion shortcuts for the given partition $\{G_{i}\}_{i=1}^{N}$ and aggregating information over the graphs $G_{i}\cup H_{i}$ in $\tilde{O}(\alpha+\beta)$ rounds, which is $\tilde{O}(D)$ for planar graphs [13].
Face-disjoint graph. The face disjoint graph $\hat{G}$ was presented in [17] as a way for simulating aggregations on (subtrees of) the dual graph $G^{*}$ in the distributed setting. This tool allows us to overcome several challenges and succeeds to perform computations on $G^{*}$ by communicating on $G$. The obvious challenge, among others, that is solved using $\hat{G}$, is that computational entities in the input (primal) graph $G$ are vertices and not faces. Moreover, a vertex on $G$ belongs to many (possibly a linear number of) faces, however, we wish to compute aggregate functions on sets of faces of $G$ using low-congestion shortcuts. Hence, to do so efficiently we want faces to be vertex-disjoint, which is the main property that $\hat{G}$ provides, in addition, it is a planar graph with diameter $O(D)$, and we can simulate $\mathsf{CONGEST}$ algorithms efficiently on it. Then, we can define each face as its own part in the low-congestion shortcuts partition, therefore, we can compute low-congestion shortcuts and computes aggregates on faces of $G$ in $\tilde{O}(D)$ rounds.
To achieve that, $\hat{G}$ is (roughly) defined as the result of duplicating all edges of $G$ so that faces of $G$ map to distinct faces of $\hat{G}$ that are both vertex and edge disjoint. Then, in our definition of $\hat{G}$, those faces in $\hat{G}$ are connected with edges if their corresponding faces in $G$ share an edge (i.e., the dual nodes of those faces are connected by an edge in $G^{*}$). See Figure 1.
Figure 1: On the left, the primal graph $G$. On the right $\hat{G}$. Notice that each face of $G$ maps to a face (black cycle) of $\hat{G}$.
Simulation in the dual graph. Our goal is to use the face disjoint graph $\hat{G}$ to simulate a minor-aggregation algorithm on the dual graph, where the vertices of a face simulate the corresponding dual node, and each dual edge is simulated by the endpoints of the corresponding primal edge. By the above discussion, our main goal is to show how to solve the PA problem on the dual graph (i.e., the input this time is a partition $\{G^{*}_{i}\}_{i=1}^{N}$ of the dual graph into connected subgraphs).
We remark that Ghaffari and Parter [17] showed how to solve specific aggregate functions on $G^{*}$ using the graph $\hat{G}$ (aggregations on each face of $G$ and sub-tree sums on $G^{*}$). For our purposes, we need something more general, so we show how to perform general part-wise aggregations on $G^{*}$. In particular, we need to perform aggregations that take into consideration the outgoing edges of each part $G^{*}_{i}$, something that was not done in [17]. This specific task requires a small modification to $\hat{G}$ compared to the one defined in [17] (see 3.1 and LABEL:appendix:_preliminaries_hat{G} for more details).
To solve the PA problem on $G^{*}$ we exploit the structure of $\hat{G}$. We prove that a PA problem in the dual graph $G^{*}$ can be translated to a corresponding PA problem in $\hat{G}$. In the graph $\hat{G}$ there is a disjoint cycle representing each face of $G$, where two such cycles are connected if there is a dual edge between the corresponding faces. Hence, if we take a partition $\{G^{*}_{i}\}_{i=1}^{N}$ of the dual graph into connected subgraphs, we can convert it to a partition of the vertices of $\hat{G}$, where each dual node $f\in G^{*}_{i}$ is replaced by the vertices of the cycle corresponding to $f$ in $\hat{G}$. Since cycles corresponding to different faces are vertex-disjoint in $\hat{G}$ we indeed get a partition.
In addition, by the properties of $\hat{G}$, it is a planar graph with diameter $O(D)$, and we can simulate $\mathsf{CONGEST}$ algorithms efficiently on $\hat{G}$. Based on this we can solve the PA problem in $\tilde{O}(D)$ rounds in $\hat{G}$, which leads to solving the PA problem on the dual graph $G^{*}$ in $\tilde{O}(D)$ rounds. For full details see Section 4. We also prove that we can simulate an extended version of the minor-aggregation model defined in [18] that allows adding $\tilde{O}(1)$ virtual nodes to the network, which is useful for our applications.
Dealing with parallel edges in $G^{*}$. The the dual graph $G^{*}$ might be a multi-graph. E.g. two faces that share two edges in $G$ correspond to two nodes that are connected with two parallel edges in $G^{*}$. For some of our applications it is useful to think about $G^{*}$ as a simple graph. For example, when computing shortest paths we would like to keep only the edge of minimum weight connecting two dual nodes, and when computing a minimum cut we want to replace multiple parallel edges by one edge with the sum of weights. Note that if $G^{*}$ was the network of communication it was trivial to replace multiple parallel edges by one edge locally, but in our case each edge on a face $f$ is known by a different vertex of $f$. While we can compute an aggregate operator of all the edges adjacent to $f$ efficiently, if a face $f$ has many different neighboring faces $g$, we need to compute an aggregate operator for each such neighboring face of $f$, which is too expensive. To overcome this challenge, we use the low arboricity  [^5] The arboricity of a graph is the minimal number of forests into which its edges can be partitioned of planar graphs to compute a low out-degree orientation of the edges of the dual graph. This orientation guarantees that from each face $f$ there are outgoing edges only to $O(1)$ other faces $g$, hence we can allow to compute an aggregate function on the edges between each such pair efficiently (each face takes responsibility to compute $O(1)$ aggregate operations). The algorithm for computing the orientation follow an algorithm of [1] that can be viewed as solving a series of $\tilde{O}(1)$ part-wise aggregation problems.
Applications. By simulating minor-aggregation algorithms on the dual graph $G^{*}$ we compute the exact minimum weighted cut in the dual graph, which by duality provides a solution to the weighted girth problem in the primal graph $G$. This approach allows us to find the weight of the minimum weight cycle, we also show that we can extend it to find the edges of the cycle, for details see Section 4. We also use this model to simulate the recent approximate SSSP by [43] on the dual graph, leading to our approximate max $st$-flow algorithm on the primal graph when $s$ and $t$ are on the same face in the given planar embedding (see Section 6.1). Since currently there are fast SSSP minor-aggregation algorithms only for undirected graphs, this approach leads to an approximate max $st$-flow algorithm in undirected planar graphs. To solve the more general version of the max $st$-flow problem, we need additional tools described next.